# Smart Contract Fuzzing Towards Profitable Vulnerabilities

### v1

@@ 4.3. RQ3 Real-World Application
Listing 4: The vulnerable burn implementation of VULN1.
Project VULN1    The project VULN1 contains a vulnerability that anyone can burn other users’ tokens as shown in Listing 4. This vulnerability is trivially caught and exploited by Verite due to matching "Unconditional Token Burn" as defined in Section 3.3. Verite can gain and maximize profit by burning tokens of the pools and manipulating the price.
Project VULN2    The project VULN2 consists of several components, and there are two contracts related to the vulnerability, Vault and Router as shown in Listing 5 and 6 respectively. Vault acts like a common token vault: receives tokens from users and mints shares in return, except supporting investing in multiple token types as shown in line 3. Similar to Uniswap, the interfaces of Vault are not intended for direct usage, and thus Router wraps the call to Vault and handles token transfers afterward.
1function mint(address sender, address token, ... unit amount) {
@@
3 uint shares = _xxxx[token];
@@
12 _xxxx[token] = shares + out;
@@
Listing 5: Vault::mint snippet of VULN2
1function xxxxx(unit amount) {
2 // Some checks ...
3 Vault.mint(msg.sender, _xxx_token, ..., amount );
4
5 _xxx_token.transferFrom(msg.sender, Vault, amount);
@@
Listing 6: Router::xxxxx snippet of VULN2
The vulnerability roots in the implementation of Vault::mint shown in Listing 5. Assume Vault is empty, and the first user tries to deposit 10 tokens by calling Router::xxxxx. An attacker could front-run his transaction with two transactions.
  • Calling Router::xxxxx with amount=1: Because Vault is empty, the condition at line 7 is true and attacker become the sole shares holder of Vault.
  • Transfer 10 _xxx_token to Vault: This essentially a donation to Vault but the balance defined in line 4 will become 11.
After the attack has finished his transactions, the user will get his shares by the line 10 because the Vault is no longer empty. The shares calculated is $(10*1)/11=0$ because of the rounding, which indicates the user does not get any share from his transaction but still transfers tokens at line 5 in Listing 6. After that, the attacker can safely withdraw all 11 tokens because there are no other shareholders.
Generally, Verite can generate the transaction sequence in less than one second because the intended sequence is short enough, and the second transaction making the donation is exactly what Action A1 defined in Table 1 does. Furthermore, the profit function in this vulnerability is monotonous, which is suitable for Verite to pick the most optimal donation amount to drain all victim’s assets.
@@ 5. Discussion
Automatic Action Synthesis    While Verite has demonstrated its effectiveness and practicality, there are still limitations in terms of its current support for action extraction. These limitations may affect the effectiveness of testing a broader range of smart contract targets. Currently, Verite supports actions for common liquidation and exchange DeFi applications and concentrates on profitable vulnerabilities. Although these presumptions suit a variety of smart contracts, given the popularity of the actions, these may not apply when testing other targets with different DeFi business models. Currently, our solution enables users to define custom actions assisted by LLM in Solidity (see more details at our website (our, 2024)), obviating the need to modify the fuzzer for enhanced usability. An interesting area of future research could be to devise adaptive methods that automate this manual process. Specifically, a domain knowledge model could first learn from the project and then be used to improve the abilities of action and abnormal action. Previous work, such as  (Li et al., 2023) and  (Wu et al., 2023), has proposed automated ways to extract high-level DeFi semantics from transactions, which can be applied to Verite for better generality. We consider this an interesting direction for future research.
@@ 7. Conclusion
To facilitate future research and follow open science policy, we will release both experiment data and artifacts of Verite on https://github.com/veritefuzz/verite once public.

### v2

@@ 4.3. RQ3 Real-World Application
Listing 5: The vulnerable burn implementation of ENDERV1.
Project ENDERV1    The project ENDERV1 contains a vulnerability that anyone can burn other users’ tokens as shown in Listing 5. This vulnerability is trivially caught and exploited by Verite due to matching "Unconditional Token Burn" as defined in Section 3.3. Verite can gain and maximize profit by burning tokens of the pools and manipulating the price.
Project ETHOSX    The project ETHOSX consists of several components, and there are two contracts related to the vulnerability, Vault and Router as shown in Listing 6 and 7 respectively. Vault acts like a common token vault: receives tokens from users and mints shares in return, except supporting investing in multiple token types as shown in line 3. Similar to Uniswap, the interfaces of Vault are not intended for direct usage, and thus Router wraps the call to Vault and handles token transfers afterward.
1function mint(address sender, address token, ... uint256 amount) {
@@
3 uint shares = _shares[token];
@@
12 _shares[token] = shares + out;
@@
Listing 6: Simplified Vault::mint snippet of ETHOSX
1function buyLongToken(uint256 amount) {
2 // ...
3 _vault.mint(msg.sender, _token, ..., amount);
4 // ...
5 _token.transferFrom(msg.sender, _vault, amount);
@@
Listing 7: Simplified Router::buyLongToken snippet of ETHOSX
The vulnerability roots in the implementation of Vault::mint shown in Listing 6. Assume Vault is empty, and the first user tries to deposit 10 tokens by calling Router::buyLongToken. An attacker could front-run his transaction with two transactions.
  • Calling Router::buyLongToken with amount=1: Because Vault is empty, the condition at line 7 is true and attacker become the sole shares holder of Vault.
  • Transfer 10 _token to Vault: This essentially a donation to Vault but the balance defined in line 4 will become 11.
After the attack has finished his transactions, the user will get his shares by the line 10 because the Vault is no longer empty. The shares calculated is $(10*1)/11=0$ because of the rounding, which indicates the user does not get any share from his transaction but still transfers tokens at line 5 in Listing 7. After that, the attacker can safely withdraw all 11 tokens because there are no other shareholders.
Generally, Verite can generate the transaction sequence in a few seconds because the intended sequence is short enough, and the second transaction making the donation is exactly what Action A1 defined in Table 1 does. Furthermore, the profit function in this vulnerability is monotonous, which is suitable for Verite to pick the most optimal donation amount to drain all victim’s assets.
@@ 5. Discussion
Automatic Action Synthesis    While Verite has demonstrated its effectiveness and practicality, there are still limitations in terms of its current support for action extraction. These limitations may affect the effectiveness of testing a broader range of smart contract targets. Currently, Verite supports actions for common liquidation and exchange DeFi applications and concentrates on profitable vulnerabilities. Although these presumptions suit a variety of smart contracts, given the popularity of the actions, these may not apply when testing other targets with different DeFi business models. Currently, our solution enables users to define custom actions assisted by LLM in Solidity (see more details at our website (our, 2025)), obviating the need to modify the fuzzer for enhanced usability. An interesting area of future research could be to devise adaptive methods that automate this manual process. Specifically, a domain knowledge model could first learn from the project and then be used to improve the abilities of action and abnormal action. Previous work, such as  (Li et al., 2023) and  (Wu et al., 2023), has proposed automated ways to extract high-level DeFi semantics from transactions, which can be applied to Verite for better generality. We consider this an interesting direction for future research.
Integrate Static Application Security Testing (SAST) Tools SAST tools like slither (Feist et al., 2023), mythril (Concensys, 2025) and securify (Tsankov et al., 2018) have been proven effective for auditing solidity projects in both industry and academia. SAST tools usually do not require complete deployment for a concrete execution environment, and thus are flexible and scalable to capture a wide range of smart contract vulnerabilities. However, Verite offers unique advantages compared to SAST tools: ① Fuzzers can provide concrete reproduction of the vulnerabilities and Verite further ensures the vulnerabilities lead to asset loss without false positives. SAST tools are usually driven by general rules and thus prone to produce false positives. ② Verite offers unique gradient based exploitation strategy, while SAST tools are usually designed to reveal but not exploit the vulnerabilities. In addition, SAST tools are complementary rather than exclusive to Verite. For instance, SAST tools can help identify potential fuzzing targets and understand the inter-action dependencies. It is also possible to enhance the proof-of-profit candidate criteria for Verite by SAST tools. We leave the strategic combination of the SAST tools with Verite as future directions.
@@ 7. Conclusion
Internal Threats to Validity
The dataset, baselines, and experimental setup in our work could introduce bias, affecting the soundness of our results. Although the dataset was constructed using all publicly known attacks on Uniswap and ERC20 tokens exploited on ETH or BSC over the past two years from DeFiHackLabs (def, 2024a), biases could arise from missing or underreported incidents and the overrepresentation of highly publicized attack types. For baselines, we reviewed all known fuzzers from the literature and either included comparison results or explained why comparisons were not feasible. In our experimental setup, we adhered to best practices from prior work (Klees et al., 2018) to reduce randomness in fuzzing evaluations.
@@
To facilitate future research and follow open science policy, we will release both experiment data and supplemental materials of Verite on  (our, 2025) and  (Kong et al., 2025) for archive purpose.
Acknowlgement
We gratefully acknowledge the support of this research by the National Research Foundation, Singapore, and the Cyber Security Agency under the National Cybersecurity R&D Programme (NCRP25-P04-TAICeN). We also thank the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-008), as well as the National Research Foundation, Prime Minister’s Office, Singapore under the Campus for Research Excellence and Technological Enterprise (CREATE) programme. Additionally, we acknowledge the support from the National Key Research and Development Program of China (2022YFB2703503). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of funding agencies.